is the solution of the equation 3x – 2 =7.
                              is the solution of the equation 3x – 2 =7.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Consider the given equation, 3x–2 = 7
Transpose -2 to RHS then it becomes 2.
3x = 7 + 2
3x = 9
x =9/3
x =3
x = 3 is the solution of the equation 3x–2 =7.

Thanks